# Clone Deterministic 3D Worlds with Geometrically-Regularized World Models

### v1

@@ 3.3 Temporal Contrastive Regularization
The output of the encoder, a sequence of latent vectors $\mathbf{z}\in\mathbb{R}^{B\times L\times\dots}$, is first passed through a projection layer (a linear layer) to produce embeddings $\mathbf{p}\in\mathbb{R}^{B\times L\times D}$  (Chen et al., 2020). These embeddings are then $L_{2}$-normalized to lie on the unit hypersphere: $\mathbf{p}^{\prime}=\frac{\mathbf{p}}{\|\mathbf{p}\|_{2}}$ (Wang & Isola, 2020). Our regularization objectives are applied to these normalized embeddings, $\mathbf{p}^{\prime}$.
Temporal Slowness Loss (${\mathcal{L}}_{\text{slow}}$). The idea of temporal slowness is that consecutive or nearby states in a trajectory should have similar latent representations, reflecting the intuition that the underlying state of the environment evolves gradually over time (Wiskott & Sejnowski, 2002). Our loss encourages all pairs of frames within the same trajectory’s context window to be close to one another on the hypersphere. This enforces that the entire trajectory segment is mapped to a compact and continuous path in the representation space, ensuring that the latent representation evolves slowly and smoothly over time. We formalize this by minimizing the average L2 distance between all pairs of embeddings within a trajectory:
@@
Latent Uniformity Loss (${\mathcal{L}}_{\text{uniform}}$). Slowness alone can lead to feature collapse (i.e., the model maps many inputs to a small region of the latent space). The uniformity loss mitigates this issue by encouraging embeddings to distribute evenly on the hypersphere. It is formally expressed as:
@@
Overall Training Objective. The complete autoencoder is trained end-to-end by minimizing a total objective function that combines the reconstruction loss, a KL-divergence term from the VAE framework, and our two proposed regularization terms. The final loss is:
@@
3.4 Practical Considerations
GRWM is designed to close the performance gap to an oracle model in a simple and general way, without requiring supervised data. Our geometric regularization module embodies this principle. Its implementation is minimalist, requiring only two additions to a standard VAE framework: a lightweight projection head and our two supplementary loss terms. This design makes GRWM a plug-and-play component that can be applied to enhance any latent generative model without needing to alter its core architecture. The computational overhead is minimal and scales naturally with batch size and trajectory length, without complex sampling or augmentation strategies. This simplicity and generality make it a practical tool for improving representation quality, turning any standard autoencoder into a powerful foundation for high-fidelity world modeling.
@@ 4.1 Setup
Baselines. Our world model framework is flexible regarding the choice of dynamics model. For a fair comparison, we select three popular dynamics models: Standard Diffusion (SD) (Alonso et al., 2024), Video Diffusion Models (VD) (Ho et al., 2022), and Diffusion Forcing (DF)  (Chen et al., 2024) [^1] If the original paper does not support actions explicitly, we treat the action as an additional condition and feed it into the network.. For each of these dynamics models, we train and compare two versions: a vanilla version that uses a standard VAE, and GRWM version. This direct comparison allows us to isolate the contribution of representation learning and assess its impact across different dynamics architectures. A detailed description of training configurations is provided in Appendix D.
Metrics. We evaluate prediction fidelity using frame-wise MSE. Since the environments are deterministic, a unique ground-truth future exists for any given initial state and action sequence. We can therefore compute the mean squared error between the predicted and ground-truth observations at each timestep in pixel space. This metric reflects the step-by-step discrepancy between predicted and actual observations across the entire trajectory. A lower frame-wise MSE indicates higher fidelity in cloning the environment’s dynamics over time.
@@ 4.2 Rollout Fidelity
To quantitatively evaluate rollout prediction fidelity, we present our main results in Figure 3. The results demonstrate a consistent advantage for our method, and the performance gap widens with rollout length. Our method (solid lines) maintains a significantly lower prediction error compared to the baseline using a vanilla VAE (dashed lines). Baseline models accumulate error quickly, causing trajectories to diverge, whereas our method maintains a much flatter error curve. This shows stronger long-term temporal consistency.
@@
These findings confirm our central hypothesis: by learning a more structured and aligned latent space, our method provides a stronger foundation for the dynamics model, leading to substantial improvements in high-fidelity prediction.
4.3 Qualitative Analysis of Long-Horizon Generation
To further probe the long-term stability, we conduct an extreme long-horizon generation task. For this experiment, we specifically select our best-performing dynamics model, Diffusion Forcing [^2] In the following discussion, we refer to the VAE combined with diffusion as VAE-WM, and our method with Diffusion Forcing as GRWM., and the Maze 9x9-CE dataset. This environment is particularly well-suited for this analysis due to its high degree of perceptual aliasing—it contains many corridors and rooms that are visually similar, making it difficult for a model to distinguish between different states based on a single frame. We use a sequence of randomly sampled actions to simulate an exploratory trajectory.
We evaluate our model on trajectories of varying lengths. To test long-horizon stability, we tasked the model with producing 10,000 consecutive frames from a single starting point. We visualize these long trajectories in Figure 4 and Figure 6. For the more complex MC-DET environment, we specifically evaluate a 100-frame rollout, as shown in Figure 5. Additional results under different starting conditions are provided in the Appendix C.
@@ 4.3 Qualitative Analysis of Long-Horizon Generation
In contrast, our GRWM generates a far more coherent and diverse trajectory because its representation is regularized to be consistent with the true underlying state manifold. As shown in Figure 4, GRWM maintains high fidelity at both 100 and 400 frames, while the VAE-WM fails. In the longer rollout (Figure 6), the sequence of frames from GRWM clearly shows movement and exploration. The appearance of the yellow wall in later frames is not a random plausible image; it is evidence of a continuous traversal through a latent space that mirrors a physically possible path in the environment. Our geometric regularization forces the model to respect the world’s structure, preventing the state-skipping and teleportation artifacts that dominate the VAE rollouts.
@@ 4.4 Latent Representation Analysis
Figure 7: Visualization of latent space structure through clustering analysis. We perform k-means clustering ($k=20$) on the latent representations of frames. Each point in the plots corresponds to a frame, positioned according to its true ($x,y$) coordinates in the environment. Points are colored based on their assigned latent cluster ID. The top row (VAE-WM) shows scattered, noisy clusters, indicating that spatially distant frames are incorrectly grouped together. The bottom row (GRWM) shows well-defined, spatially coherent clusters, demonstrating that our learned latent space is structurally aligned with the environment’s true state manifold.
@@ 5 Ablation Studies
We conduct ablation studies to validate the contribution of our core components and design choices. Specifically, we examine four aspects: (1) the necessity of the two core loss terms, (2) the role of the projection head, (3) the impact of latent dimension, and (4) the effect of critical design choices on model performance. The detailed results are provided in Appendix B.
6 Discussion and Conclusion
In this work, we addressed the challenge of high-fidelity cloning for deterministic, closed environments. We began from a simple yet powerful premise: representation matters. While much of the field has focused on developing more powerful dynamics models, our work makes a definitive case that the primary bottleneck to long-horizon prediction is the quality of the latent space in which those dynamics operate. We introduced Geometrically-Regularized World Models (GRWM), a framework that learns a latent space structurally aligned with the environment’s true state manifold through a combination of a temporal-contextual architecture and a novel geometric regularization loss. Our experiments consistently demonstrated that by improving the representation, we could systematically enhance the performance of various dynamics models, significantly reducing error accumulation and preventing the catastrophic trajectory divergence that plagues standard methods.
Limitations and Future Vision.
While our geometric regularization substantially improves latent space topology, it alone does not guarantee that the learned representation is equivalent to the true, irreducible physical state. Any residual misalignment, however small, could eventually compound over extremely long horizons. This limitation points toward the ultimate goal and grand challenge for the field: the unsupervised discovery of the true underlying generative factors of an environment.
An optimal representation should capture these physical states, reducing high-dimensional sensory signals to their simplest, most informative form without losing information critical for prediction. Recent advances in self-supervised learning, vector quantization, and learned compression have made progress toward this goal, but the challenge remains open. The path forward lies in developing methods that can not only learn a geometrically sound manifold, as we have done, but can also automatically disentangle the true factors of variation. A model equipped with such a perfect representation would not just be cloning an environment; it would be understanding its fundamental laws.
In conclusion, our work repositions the problem of high-fidelity world modeling as one of representation learning first, and dynamics modeling second. By shifting the focus from the complexity of the transition function to the geometry of the state space, we have taken a significant step toward building robust, long-horizon predictive models.
Appendix
Appendix A LLM Usage
LLMs were used only to polish language and improve writing fluency; all research content is solely by the authors.
Appendix B Ablation Studies
@@
B.1 Importance of Core Regularization Terms
@@
B.2 Role of the Projection Head
@@
B.3 Analysis on Latent Dimension
@@
B.4 Design of the Slowness Loss
@@
Appendix C Additional Rollout Visualizations
@@
Appendix D Training Details
@@
Appendix E Dataset Details
@@
For the Maze environment, we fix the random seed to generate a consistent set of maps. We use Memory-Maze Environment (Pasukonis et al., 2022). The rendered images are obtained using the MuJoCo engine. The agent has a discrete action space consisting of {move forward, turn left, turn right}. Trajectories are collected with a noisy A* algorithm to ensure sufficient coverage of the maze.
@@
For the Minecraft environment, we adopt the map from Gornet & Thomson (2024). To restrict exploration, we enclose the area with wooden fences. Data collection is performed using the Malmo framework, with the same action space as in the Maze environment. The noisy A* policy is again used to generate diverse trajectories.

### v2

@@ 5.2 Temporal Contrastive Regularization
The output of the encoder, a sequence of latent vectors $\mathbf{z}\in\mathbb{R}^{B\times L\times\dots}$, is first passed through a projection layer (a linear layer) to produce embeddings $\mathbf{p}\in\mathbb{R}^{B\times L\times D}$ [^4] BB denotes the batch size, $L$ the sequence length, and $D$ the embedding dimension. (chen2020simple). These embeddings are then $L_{2}$-normalized to lie on the unit hypersphere: $\mathbf{p}^{\prime}=\frac{\mathbf{p}}{\|\mathbf{p}\|_{2}}$ (wang2020understanding). Our regularization objectives are applied to these normalized embeddings, $\mathbf{p}^{\prime}$.
Temporal Slowness Loss (${\mathcal{L}}_{\text{slow}}$).
The idea of temporal slowness is that consecutive or nearby states in a trajectory should have similar latent representations, reflecting the intuition that the underlying state of the environment evolves gradually over time (wiskott2002slow). Our loss encourages all pairs of frames within the same trajectory’s context window to be close to one another on the hypersphere. This enforces that the entire trajectory segment is mapped to a compact and continuous path in the representation space, ensuring that the latent representation evolves slowly and smoothly over time. We formalize this by minimizing the average L2 distance between all pairs of embeddings within a trajectory:
@@
Latent Uniformity Loss (${\mathcal{L}}_{\text{uniform}}$).
Slowness alone can lead to feature collapse (i.e., the model maps many inputs to a small region of the latent space). The uniformity loss mitigates this issue by encouraging embeddings to distribute evenly on the hypersphere. It is formally expressed as:
@@
Overall Training Objective.
The complete autoencoder is trained end-to-end by minimizing a total objective function that combines the reconstruction loss, a KL-divergence term from the VAE framework, and our two proposed regularization terms. The final loss is:
@@ 6.1 Setup
Baselines.
State-of-the-art latent generative world models generally adopt a standard VAE as the backbone for representation learning. Our method does not modify the dynamics model itself. Therefore, when comparing with existing sota approaches, we pair our representation module with the same dynamics models they use. We select three sota latent generative world models: Standard Diffusion (SD) (alonso2024diffusionworldmodelingvisual), Video Diffusion Models (VD) (ho2022video), and Diffusion Forcing (DF)  (chen2024diffusion). A detailed description of training configurations is provided in Appendix B.
Metrics.
We evaluate prediction fidelity using frame-wise MSE. Since the environments are deterministic, a unique ground-truth future exists for any given initial state and action sequence. We can therefore compute the mean squared error between the predicted and ground-truth observations at each timestep in pixel space. This metric reflects the step-by-step discrepancy between predicted and actual observations across the entire trajectory. A lower frame-wise MSE indicates higher fidelity in cloning the environment’s dynamics over time.
@@ 6.2 Rollout Fidelity
To quantitatively evaluate rollout prediction fidelity, we present our main results in Figure 3. The results demonstrate a consistent advantage for GRWM, and the performance gap widens with rollout length. GRWM (solid lines) maintains a significantly lower prediction error compared to the baseline using a vanilla VAE (dashed lines). Baseline models accumulate error quickly, causing trajectories to diverge, whereas our method maintains a much flatter error curve. This shows stronger long-term temporal consistency.
@@
6.3 Qualitative Long-Horizon Results
To further probe the long-term stability, we conduct an extreme long-horizon generation task. For this experiment, we specifically select sota world model, Diffusion Forcing [^5] In the following discussion, we refer to the original Diffusion Forcing as VAE-WM, and GRWM with Diffusion Forcing as GRWM., and the M9x9-DET and MC-DET dataset. We use a sequence of randomly sampled actions to simulate an trajectory.
We evaluate our model on trajectories of varying lengths. To test long-horizon stability, we tasked the model with producing 10,000 consecutive frames from a single starting point. We visualize these long trajectories in Figure 4 and Figure 6. For the more complex MC-DET environment, we specifically evaluate a 100-frame rollout, as shown in Figure 5. Additional results under different starting conditions are provided in the Appendix A.
@@ 6.3 Qualitative Long-Horizon Results
In contrast, GRWM generates a far more coherent and diverse trajectory because its representation is regularized to be consistent with the true underlying state manifold. As shown in Figure 4, GRWM maintains high fidelity at both 100 and 400 frames, while the VAE-WM fails. In the longer rollout (Figure 6), the sequence of frames from GRWM clearly shows movement and exploration. The appearance of the yellow wall in later frames is not a random plausible image; it is evidence of a continuous traversal through a latent space that mirrors a physically possible path in the environment. Our geometric regularization forces the model to respect the world’s structure, preventing the state-skipping and teleportation artifacts that dominate the VAE rollouts.
@@ 6.4 Latent Representation Analysis
Figure 7: Visualization of latent space structure through clustering analysis. We perform $k$-means clustering ($k=20$) on the latent representations of frames. Each point in the plots corresponds to a frame, positioned according to its true ($x,y$) coordinates in the environment. The $(x,y)$ coordinates are normalized and lie within [1,1][-1,1]. Points are colored based on their assigned latent cluster ID. The top row (VAE-WM) shows scattered, noisy clusters, indicating that spatially distant frames are incorrectly grouped together. The bottom row (GRWM) shows well-defined, spatially coherent clusters, demonstrating that our learned latent space is structurally aligned with the environment’s true state manifold.
@@ 7 Ablation Studies
7.1 Importance of Core Regularization Terms
@@
7.2 Role of the Projection Head
@@
7.3 Analysis on Latent Dimension
@@
7.4 Design of the Slowness Loss
@@
8 Conclusion
In this work, we studied the problem of building high-fidelity world models for deterministic, closed 3D environments. Our starting point is that the representation quality is crucial. Much prior work has focused on designing stronger dynamics models, but our results show that the main limitation in long-horizon prediction comes from the structure of the latent space in which the dynamics operate in. We apply temporal contrastive principles to shape the latent space in latent generative world model. This allows the dynamics models to reach their full capability. It narrows the gap to the oracle and supports our main hypothesis: representation quality is a key factor for achieving reliable long-horizon prediction.
Appendix
Appendix A Additional Rollout Visualizations
@@
Appendix B Training Details
@@
Appendix C Dataset Details
@@
For the Maze environment, we fix the random seed to generate a consistent set of maps. We use Memory-Maze Environment (pasukonis2022memmaze). The rendered images are obtained using the MuJoCo engine. The agent has a discrete action space consisting of {move forward, turn left, turn right}. Trajectories are collected with a noisy A* algorithm to ensure sufficient coverage of the maze.
@@
For the Minecraft environment, we adopt the map from gornet2024automated. To restrict exploration, we enclose the area with wooden fences. Data collection is performed using the Malmo framework, with the same action space as in the Maze environment. The noisy A* policy is again used to generate diverse trajectories.